After a financial crisis hits the country of Arcadia, 22 million people become unemployed. If 5555 million individuals are lucky enough to keep their jobs, what is the unemployment rate? Please specify your answer to two decimal places.

unemployment rate:________________%

Which of the choices is most directly related to cyclical unemployment?

A.) the age distribution of the country

B.) an increased use of the Internet for job searches

C.) recessions

D.) an increase in the number of new college graduates looking for work

Which individual is out of the labor force?

A.) Betty is without a job, is able to work, and is actively looking for work.

B.) Olivia is working 8 hours per week.

C.) Matthew is working 40 hours per week.

D.) John is without a job, is able to work, but is not actively looking for work.
Answer #1

1.

Labor force = employed + unemployed = 5555 + 22 = 5577m

Unemployment rate = unemployed / labor force * 100 = 22 / 5577 * 100 = 0.3944% ~ 0.39 % (two decimal places)

2. Correct answer is option C. ie.  recessions

3. Correct answer is option D. ie. John,

because John would be classified as discouraged worker and discouraged workers are not under labor force